OpenStudy (anonymous):

solve the quadradratic equation by factoring xsquared-49=0

OpenStudy (anonymous):

(x-7)(x+7)=0 by the zero product property x-7=0 or x+7=0 so x=-7,7

OpenStudy (anonymous):

thanks you please explian?